OpenStudy (anonymous):

Find the derivative of the function. y = cot2(sin θ)

OpenStudy (anonymous):

chain rule thrice

OpenStudy (anonymous):

wait, it's cot^2

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

y=cot^2(sintheta)

OpenStudy (anonymous):

derivative of something squared is 2 times something, times the derivative of something derivative of cotangent of something is - cosecant squared of something , times the derivative of something derivative of sine is cosine

OpenStudy (anonymous):

peal the off one at a time \[2\cot(\sin(x))\times (-\csc^2(\sin(x))\times \cos(x)\]

OpenStudy (anonymous):

ohhh, i see. i was trying to combine two things in one step

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw